OpenStudy (anonymous):

How would i factor a simple question like 3a^2 + 2

jimthompson5910 (jim_thompson5910):

3a^2 and 2 have nothing in common (except 1), so that cannot be factored

OpenStudy (anonymous):

i just still dont understand....oh boy..

OpenStudy (anonymous):

what about one like 3a^2 + 9?

jimthompson5910 (jim_thompson5910):

the GCF of 3 and 2 is 1 there are no variables common between them, which is why the GCF of 3a^2 and 2 is 1

jimthompson5910 (jim_thompson5910):

3a^2 and 9 have a 3 in common though

jimthompson5910 (jim_thompson5910):

so you can factor out 3 to get 3a^2 + 9 = 3(a^2 + 3)

OpenStudy (anonymous):

Oh i see how that one works now. thanks. and for the one i just asked....what does it mean by putting it in factored form?

jimthompson5910 (jim_thompson5910):

it just means factor it as much as possible

jimthompson5910 (jim_thompson5910):

example: if you factor the number 10, then you would get 2*5

jimthompson5910 (jim_thompson5910):

2*5 is the factored form

OpenStudy (anonymous):

like how would i factor that one out...?

jimthompson5910 (jim_thompson5910):

to factor 3a^2 + 9 completely, you would get 3a^2 + 9 = 3(a^2 + 3) and that's all you can do

jimthompson5910 (jim_thompson5910):

3(a^2 + 3) is the factored form

OpenStudy (anonymous):

OHHHHHH, THANK YOU SO MUCH. you are a hero

jimthompson5910 (jim_thompson5910):

you're welcome, glad it's making sense now

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

wait wait...

OpenStudy (anonymous):

what if in the directions it said x^2 + 3x + 2

OpenStudy (anonymous):

to factor that expression? would that work the same way?

jimthompson5910 (jim_thompson5910):

sort of...but you would use a different trick now

jimthompson5910 (jim_thompson5910):

the last term is 2 the middle coefficient is 3

jimthompson5910 (jim_thompson5910):

you want to find two numbers that multiply to 2 (last term) AND add to 3 (middle coefficient) at the same time

jimthompson5910 (jim_thompson5910):

which two numbers do this?

OpenStudy (anonymous):

well... 1 x 2 = 2 and 2 + 1 = 3 so.... 2?

OpenStudy (anonymous):

i think......

jimthompson5910 (jim_thompson5910):

good, so the two numbers are 1 and 2

jimthompson5910 (jim_thompson5910):

which means x^2 + 3x + 2 factors to (x + 1)(x + 2)

jimthompson5910 (jim_thompson5910):

you can check by expanding (x + 1)(x + 2) out and you'll get x^2 + 3x + 2 back again

OpenStudy (anonymous):

oh yay! I see how that works. what if their was a number in front of the x

jimthompson5910 (jim_thompson5910):

then it would be slightly different, but you would use the same basic idea

jimthompson5910 (jim_thompson5910):

do you have anything in mind?

OpenStudy (anonymous):

as in 4n^2 - 20n +25

jimthompson5910 (jim_thompson5910):

ok first you would multiply 4 (the first coefficient) and 25 (last term) to get 4*25 = 100

jimthompson5910 (jim_thompson5910):

now you must find two numbers that a) multiply to 100 AND b) add to -20 at the same time

OpenStudy (anonymous):

uhhh

OpenStudy (anonymous):

i understand how to do it now but i cant figure out what goes into each

jimthompson5910 (jim_thompson5910):

how about -1 and -100 -1 + (-100) = -101 no that doesn't work

jimthompson5910 (jim_thompson5910):

-2 and -50 ? -2 + (-50) = -52 nope, doesn't work either

jimthompson5910 (jim_thompson5910):

-4 and -25? -4 + (-25) = -29 still no

jimthompson5910 (jim_thompson5910):

-5 and -20 ? -5 + (-20) = -25 no, but we're getting closer

jimthompson5910 (jim_thompson5910):

-10 and -10? -10 + (-10) = -20 which works

OpenStudy (anonymous):

yeah i see that

jimthompson5910 (jim_thompson5910):

so we break up the -20n into -10n-10n to get this 4n^2 - 10n - 10n +25

jimthompson5910 (jim_thompson5910):

then you would factor by grouping

OpenStudy (anonymous):

like terms?

jimthompson5910 (jim_thompson5910):

first group the terms into pairs 4n^2 - 10n - 10n +25 (4n^2 - 10n) + (-10n +25)

jimthompson5910 (jim_thompson5910):

then factor each group separately

jimthompson5910 (jim_thompson5910):

(4n^2 - 10n) + (-10n +25) 2n(2n - 5) + (-10n +25) 2n(2n - 5) - 5(2n - 5) see what to do next?

OpenStudy (anonymous):

kinda ish

jimthompson5910 (jim_thompson5910):

what do you think the next step is

OpenStudy (anonymous):

distributing

jimthompson5910 (jim_thompson5910):

yes in a way, but in reverse

jimthompson5910 (jim_thompson5910):

notice we have the common term of (2n-5)

OpenStudy (anonymous):

factoring outt.... i think

jimthompson5910 (jim_thompson5910):

so we factor that out to get this (2n - 5)(2n - 5)

OpenStudy (anonymous):

ohhh i see

jimthompson5910 (jim_thompson5910):

if we had x(2n-5) + y(2n-5) we would factor out the (2n-5) to get (x+y)(2n-5)

OpenStudy (anonymous):

oh i think i get it thank you!!!! gold star for you!

jimthompson5910 (jim_thompson5910):

ok glad it's all making sense, you'll be a pro in no time and you'll have to change your user name lol

OpenStudy (anonymous):

haha!! i just never grasp the concepts with my teacher!! haha hopefully i can change it though!!

jimthompson5910 (jim_thompson5910):

you can definitely change it, just keep practicing and you'll be fine

OpenStudy (anonymous):

will do!!!! haha

jimthompson5910 (jim_thompson5910):

lol good, and also try to get things in terms of your own language no sense in repeating the same thing over and over again if you don't know how to translate

OpenStudy (anonymous):

thats so true! i dont even know most terms! its crazy! thanks for the advice:)

jimthompson5910 (jim_thompson5910):

you're welcome, you might do well to write them on flash cards or something. That usually helps me

OpenStudy (anonymous):

yeah ill do that forsure!! and jusst wondering, are you in highschool?

jimthompson5910 (jim_thompson5910):

no I'm in college learning to become a teacher

OpenStudy (anonymous):

ohhh man! thats really cool! hope you are successful!!

jimthompson5910 (jim_thompson5910):

me too, thanks lol

OpenStudy (anonymous):

no problem!